OpenStudy (anonymous):

A periodic function f has a period of 12. If f(7) = -2 and f(11) = 9, determine the value of f(43).

OpenStudy (anonymous):

Why is the answer -2? o.O

OpenStudy (anonymous):

I am so confused D:

OpenStudy (shubhamsrg):

f(11) is of no use to this ques since period is 12,,we have f(7)=f(19)=f(31)=f(43) so the ans is -2

OpenStudy (anonymous):

ohhh, i see. Alright, thank you!